OpenStudy (anonymous):

Use the table below to answer this question:

OpenStudy (anonymous):

x | y -1| 1 1 | 5 3 | 9

OpenStudy (anonymous):

Find the average rate of change for the given function from x = 1 to x = 3. −4 −2 2 4

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Points: (1, 5), (3, 9) m = y2 - y1/x2 - x1 m = 9 - 5/3 - 1 m = 4/2 m = 2